pps For quick access, place your bookmarks here on the bookmarks bar. Import bookmarks now. Saved Help Seve&Exit Quiz 2 4 Dehner Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data: Total direct labor-hours Total fixed manufacturing overhead cost Variable manufacturing overhead per direct labor-hour $5.80 78,000 5460, 280 Recently. Job P951 was completed with the following characteristics Nunber of units in the job Total direct labor-hours Direct/ materials Direct labor cost 50 1e0 5 720 57,800 The unt oroduct cost for Job P95119s fo (Round your intermedinte calculetions to 2 decimel pleces)
Answer #1
Predetermined overhead rate = 5+(460200/78000)= $10.90
Direct materials 710
Direct labor 7800
Overhead applied 90 =100*0.9
Total manufacturing cost 8600
Units 50
Units product cost 172